Econ 2)1 Given the following figures, calculate savings, MPC, MPS, and Multiplier (20 points) 1. C S MPC MPS APC APS GDP 325 320
Answer #1

We assume that GDP is increasing continuously at 25. Hence change in Y is 25. Change in C is 20. This implies MPC = 20/25 = 0.8. Savings when Y = 300, and C = 300 so saviing = 0.

Use APC = C/Y and APS = S/Y for a given level of Y and C. MPS = 1 - MPC.

GDP C S MPC MPS APC APS
250 260 -10 0.8 0.2 1.04 -0.04
275 280 -5 0.8 0.2 1.02 -0.02
300 300 0 0.8 0.2 1.00 0.00
325 320 5 0.8 0.2 0.98 0.02
350 340 10 0.8 0.2 0.97 0.03
375 360 15 0.8 0.2 0.96 0.04